OpenStudy (anonymous):

At the Wild Thing Zoo, you can rent a motorized cart to tour the grounds for a $4 initial charge and $4 per hour. An equation that represents the total cost y for x hours is y = 4 + 4x. At Safari Zoo, you can rent the same cart for a $3 inital charge and $5 for hour. An equation that represents the total cost y for x hours is y = 3 + 5x. For how many hours is the total charge the same?

OpenStudy (anonymous):

set \[4+4x=3+5x\] and solve for \(x\) subtract \(4x\) from both sides to get \[4=3+x\] then subtract 3to get \[x=1\]

OpenStudy (anonymous):

thank you so much! it took a me a while to get it but though. thanks a bunch!

OpenStudy (anonymous):

yw